OpenStudy (anonymous):

–10 • 2 ÷ (–4) + 8 • 5

OpenStudy (calculusxy):

Use PEMDAS even though here in London we do BEDMAS, however its the same thing. Okay so lets start. -10*2/(-4) + 8*5 -20/(-4) + 40 5+40 45

OpenStudy (anonymous):

-10x2=-20 -20x-4=5 5+8x5=45